# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ Appendix C Proof of Proposition 1
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]+\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]$
@@
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]+\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]$
@@
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}+\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right].$
@@
For the second equation, we used that $P(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})=P(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},\mathbf{M})$ and that the model input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$ depends on the masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ only through their product $\mathbf{M}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$, for which the order of the masks does not play a role, thus the role of the two masks is exchangeable. For the last equation, we used that the masks $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}$ and $\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ are orthogonal.
@@
$\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}\Bigg{[}\left|\left|\big{(}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{M}+\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}+\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\big{)}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\Bigg{]}\stackrel{{\scriptstyle(*)}}{{=}}\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{\text{0}}}}\Bigg{[}\left|\left|\big{(}\mathbf{I}-\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}^{C}\big{)}\mathbf{F}\big{(}f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\Bigg{]},$
where the equality (*)(*)( * ) holds because
@@
$=\mathbf{I}-(\mathbf{I}-\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})(\mathbf{I}-\mathbf{M})$
@@
Appendix D Details on Experiments
Here, we provide additional technical details of our experiments discussed in Section 2.2. For all experiments, we used the Adam optimizer [KB15] with a constant learning rate of $4\cdot 10^{-4}$, which is the default learning rate of IsoNet. We found that DeepDeWedge is not sensitive to the choice of the learning rate. The remaining details are as follows:
D.1 Details on Section 2.2.1
  • FBP and sub-tomograms: We performed all FBP reconstructions with a ramp filter in Python using the library tomosipo [Hen+21]. After reconstruction, we downsampled all tomograms by a factor of 6 using average pooling, which resulted in a final voxel size of 13.02 Å. To these tomograms, we applied IsoNet’s CTF deconvolution routine with parameters as described by [Liu+22, ] in their paper. For model fitting, we extracted sub-tomograms of shape $80\times 80\times 80$. As wide regions of the tomograms contain only ice and no ribosomes, we used the make mask routine implemented in IsoNet to generate a mask of the non-empty regions. After extracting the sub-tomograms, we selected only those that contain at least 40% sample according to the mask. For the refinement with the fitted model, we also used sub-tomograms of shape $80\times 80\times 80$ voxels, but extracted them without masking and with an overlap of 40 voxels.
  • Details on IsoNet: We fitted IsoNet for 45 iterations (10 epochs per iteration) on 423 sub-tomograms. We used the default noise schedule for fitting, which starts adding noise with level 0.05 in iteration 11 and then increases the noise level by 0.05 every 5 iterations. We set the noise_mode parameter, which determines the distribution of the additional noise used for noisier2noise-like denoising, to noFilter, which corresponds to Gaussian noise. We also tried setting the noise mode to ramp, as the tomograms were reconstructed with FBP with a ramp filter, but this gave worse results. In their paper, [Liu+22, ] fitted IsoNet for 30 iterations using similarly many sub-tomograms and the same number of epochs. However, we found that 45 iterations gave visually more appealing results compared to the images shown in the IsoNet paper.

### v3

@@ Appendix 3 Proof of Proposition 1
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]+\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]$
@@
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]+\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right]$
@@
$=\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}}\left[\left|\left|(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}+\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\right].$
@@
For the second equation, we used that $\mathrm{P}(\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}})=\mathrm{P}(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},\mathbf{M})$ and that the model input ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}}$ depends on the masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ only through their product $\mathbf{M}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$, for which the order of the masks does not play a role, thus the role of the two masks is exchangeable. For the last equation, we used that the masks $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}$ and $\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ are orthogonal.
@@
$\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}\Bigg{[}\left|\left|\big{(}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{M}+\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}+\mathbf{M}^{C}\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\big{)}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\Bigg{]}\stackrel{{\scriptstyle(*)}}{{=}}\mathop{\mathbb{E}}_{\mathbf{M},\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{\text{0}}}}\Bigg{[}\left|\left|\big{(}\mathbf{I}-\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}^{C}\mathbf{M}^{C}\big{)}\mathbf{F}\big{(}\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}^{\text{0}}})-\mathbf{v}^{*}\big{)}\right|\right|_{2}^{2}\Bigg{]},$
where the equality ()(*)( ∗ ) holds because
@@
Appendix 4 Details on Main Experiments
Here, we describe the technical details of all the experiments presented in the main paper.
4.1 Experiments on Purified S. Cerevisiae 80S Ribosomes
The EMPIAR-10045 dataset contains 7 tilt series collected from samples of purified S. Cerevisiae 80S Ribosomes. All tilt series are aligned and consist of 41 projections collected at tilt angles from $-60^{\circ}$ to $+60^{\circ}$ with a $3^{\circ}$ increment.
  • FBP and sub-tomograms: We performed all FBP reconstructions with a ramp filter in Python using the library tomosipo [31]. After reconstruction, we downsampled all tomograms by a factor of 6 using average pooling, which resulted in a final voxel size of 13.02 Å. To these tomograms, we applied IsoNet’s CTF deconvolution routine with parameters as described by Liu et al. [4] in their paper. For model fitting, we extracted sub-tomograms of shape $80\times 80\times 80$. As wide regions of the tomograms contain only ice and no ribosomes, we used IsoNet’s mask generation tool with the default parameters to produce a mask of the non-empty regions. After extracting the sub-tomograms, we selected only those that contain at least 40% sample according to the mask. For the refinement with the fitted model, we also used sub-tomograms of shape $80\times 80\times 80$ voxels but extracted them without masking and with an overlap of 40 voxels.
  • Details on IsoNet: We fitted IsoNet for 45 iterations (10 epochs per iteration) on 423 sub-tomograms. We used the default noise schedule for fitting, which starts adding noise with level 0.05 in iteration 11 and then increases the noise level by 0.05 every 5 iterations. We set the noise_mode parameter, which determines the distribution of the additional noise used for Noisier2Noise-like denoising, to noFilter, which corresponds to Gaussian noise. We also tried setting the noise mode to ramp, as the tomograms were reconstructed with FBP with a ramp filter, but this gave worse results. In their paper, Liu et al. [4] fitted IsoNet for 30 iterations using similarly many sub-tomograms and the same number of epochs. However, we found that 45 iterations gave visually more appealing results compared to the images shown in the IsoNet paper.